Good day,

The reaction 2CH4(g)⇌C2H2(g)+3H2(g) has an equilibrium constant of K = 0.154.
If 6.60 mol of CH4, 4.35 mol of C2H2, and 11.65 mol of H2 are added to a reaction vessel with a volume of 6.00 L , what net reaction will occur?
A)The reaction will proceed to the left to establish equilibrium.
B)The reaction will proceed to the right to establish equilibrium.
C)No further reaction will occur because the reaction is at equilibrium.

Thanks
1
Expert's answer
2020-08-09T08:43:09-0400

At first, concentrations should be found: C=n/V.

C(CH4)=6.6/6=1.1(mol/L);

C(C2H2)=4.35/6=0.725(mol/L);

C(H2)=11.65/6=1.94(mol/L).

Then the reaction quotient should be defined:

Q=[C2H2]×[H2]3/[CH4]2

Q=0.725×1.943/1.12=4.37.

Finally, reaction quotient should be compared with equilibrium constant:

Q>K, so the reaction will proceed to the left.

Answer: A.